OpenStudy (anonymous):

The legs of a right triangle are 10 centimeters, & 24 centimeters long. what is the length of the hypotenuse A. 22 centimeters B.26 centimeters C.30 centimeters D. 34 centimeters

OpenStudy (anonymous):

Use pythagoream theorem. a^2 + b^2 = c^2 10^2 + 24^2 = c^2 100 + 576 = c^2 676 = c^2 Square root both sides c = 26 centimetres.